Refer to the von-Thünen model where a city’s market reaches to the left and to the right. The firm is competitive and its price equals marginal cost. The firm’s marginal cost is c=1; the market reach is x=14 to the left and x=14 to the right. Marginal transportation cost is 6.

(a) What is customers’ WTP per unit of the good in question?

(b) What is the rent for a person located at x=10?
Answer #1

a)

WTP = c+tX

C= marginal cost, t = marginal transportation cost , x=14

WTP = 1+6*14 = 85

B)

Here X= 10, t =6,c=1

Rent = WTP-(c+tX)

Rent = 85-61 =24